40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 10 15:41:37 2019, aaron, Update, IOO, WFS measurements Screenshot_from_2019-09-10_18-51-28.pngmctrend_190910.png
    Reply  Wed Sep 11 10:26:56 2019, aaron, Update, IOO, WFS measurements 
       Reply  Wed Sep 11 14:37:43 2019, aaron, Update, IOO, WFS measurements 20190911_WFS.jpg20190911_WFS_2.jpg
          Reply  Thu Sep 12 12:42:31 2019, aaron, Update, IOO, WFS measurements WFS2_TFs.pdfD755499D-9FDF-4E2B-BFC1-016B459DD35D.jpeg
             Reply  Fri Sep 13 10:36:03 2019, aaron, Update, IOO, WFS measurements TF-.pngWFS2_TFs.pdf
                Reply  Mon Sep 16 12:38:59 2019, aaron, Update, IOO, WFS measurements 
                Reply  Thu Oct 3 11:38:35 2019, aaron, Update, IOO, WFS measurements 
          Reply  Mon Sep 30 13:20:55 2019, aaron, Update, IOO, shot noise measurement 
       Reply  Fri Sep 13 10:53:40 2019, aaron, Update, IOO, WFS loop measurements 
          Reply  Mon Sep 16 05:08:04 2019, rana, Update, IOO, WFS loop measurements LSCsensingMatrix.ipynbASCsensingMatrix.ipynb
             Reply  Mon Sep 16 11:55:58 2019, rika, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 09:41:48 2019, gautam, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 10:47:44 2019, rika, Update, IOO, WFS loop measurements MC2.pdf
                   Reply  Wed Sep 18 14:45:52 2019, rika, Update, IOO, WFS loop measurements Screenshot_from_2019-09-18_18-15-34.png
             Reply  Tue Sep 17 10:34:48 2019, aaron, Update, IOO, WFS loop measurements 
                Reply  Tue Oct 8 20:39:42 2019, aaron, Update, IOO, WFS loop measurements 
                   Reply  Wed Oct 9 09:37:28 2019, aaron, Update, IOO, WFS loop measurements 
Message ID: 14874     Entry time: Thu Sep 12 12:42:31 2019     In reply to: 14872     Reply to this: 14875
Author: aaron 
Type: Update 
Category: IOO 
Subject: WFS measurements 

[rika, aaron]

At Seiji and Gautam's suggestion, we added an additional RF photodiode (NewFocus 1611) to the system so we can calibrate our transfer functions. The configuration is now laser -> BS --> lenses -> QPD and BS --> lenses -> RFPD. We added lenses to get the beams focused on the RFPD and QPD heads, and are again set up for TF measurement.

We took the following data. These parameters were consistent across all measurements:

  • 1kHz IF BW
  • log sweep with 801 points
  • 32 averages
  • auto attenuation
  • -10 dBm excitation amplitude
  • 19.2 mA DC current to the laser
  • The DC level of the reference PD is -, and with the beam blocked (dark current) it is
Measurement file parameters
WFS2_SEG1 / RFPD
TFAG4395A_12-09-2019_155901.txt
100 MHz - 500 MHz
WFS2_SEG1 / RFPD TFAG4395A_12-09-2019_160811.txt 10 MHz - 100 MHz
WFS2_SEG1 / RFPD
TFAG4395A_12-09-2019_170234.txt
100 kHz - 10 MHz
WFS2_SEG2 / RFPD AG4395A_12-09-2019_183125.txt 100 MHz - 500 MHz
WFS2_SEG2 / RFPD TFAG4395A_12-09-2019_183614.txt 10 MHz - 100 MHz
WFS2_SEG2 / RFPD TFAG4395A_12-09-2019_183930.txt 100 kHz - 10 MHz
WFS2_SEG3 / RFPD TFAG4395A_12-09-2019_225243.txt 100 MHz - 500 MHz
WFS2_SEG3 / RFPD TFAG4395A_12-09-2019_225601.txt 10 MHz - 100 MHz
WFS2_SEG3 / RFPD TFAG4395A_12-09-2019_225922.txt 100 kHz - 10 MHz
WFS2_SEG4 / RFPD
TFAG4395A_12-09-2019_230758.txt
100 MHz - 500 MHz
WFS2_SEG4 / RFPD TFAG4395A_12-09-2019_232058.txt 10 MHz - 100 MHz
WFS2_SEG4 / RFPD TFAG4395A_12-09-2019_234447.txt 100 kHz - 10 MHz

After taking the data for segment 1, I moved the beam to segment 2. The beam didn't fit on segment 2 without partially illuminating segment 1 (tested by maximizing the signal on segment 2, then blocking the beam. If the beam is entirely on one segment, only that segment should be effected; in this case, we found that segment 1's DC signal also changed when the beam was blocked). We readjusted the telescoping lenses to get the beam a bit smaller, and now the beam fits on segment 2. We know it is entirely on segment 2 because small beam movements do not change the signal on segment 2.

We are trying to take the remaining data, but AGmeasure keeps hanging while sending the data (after taking the measurement, over 10 min). We tried restarting the network analyzer to no avail. I was able to grab the data by cancelling the measurement and running

AGmeasure --getdata -i vanna

I've uploaded the spectrum for segment 1 in the meantime. Zero model is on the way.

When I finished up the measurements on WFS2, I removed the cables from the AP table and closed the cover.

EDIT: I forgot to switch the LEMO connector to measure the other segments, so we measured the RF signal from segment 1 even when the beam was on segments 2-4. We'll have to try again tomorrow.

Attachment 1: WFS2_TFs.pdf  30 kB  Uploaded Thu Sep 12 19:10:24 2019  | Hide | Hide all | Show all
WFS2_TFs.pdf
Attachment 2: D755499D-9FDF-4E2B-BFC1-016B459DD35D.jpeg  3.364 MB  Uploaded Mon Sep 16 18:58:20 2019  | Show | Hide all | Show all
ELOG V3.1.3-